### 23 Jan 2024 v3

# On the Existence of Steady-State Solutions to the Equations Governing Fluid Flow in Networks

Abstract
The steady-state solution of fluid flow in pipeline infrastructure networks driven by junction/node potentials is a crucial ingredient in various decision support tools for system design and operation. While the non-linear system is known to have a unique solution (when one exists), the absence of a definite result on existence of solutions hobbles the development of computational algorithms, for it is not possible to distinguish between algorithm failure and non-existence of a solution. In this letter we show that a unique solution exists for such non-linear systems if the term solution is interpreted in terms of potentials and flows rather than pressures and flows. The existence result for flow of natural gas in networks also applies to other fluid flow networks such as water distribution networks or networks that transport carbon dioxide in carbon capture and sequestration. Most importantly, by giving a complete answer to the question of existence of solutions, our result enables correct diagnosis of algorithmic failure, problem stiffness and non-convergence in computational algorithms.
{IEEEkeywords}
[…]
1 Introduction
[…]
FLUID flow in pipeline infrastructure networks driven by junction/node potentials through a network consisting of pipelines and pressure regulating devices (like pumps/compressors) is relevant to energy distribution and environmental remediation, such as in delivery of hydrogen/natural gas, water distribution, or transport of carbon dioxide in carbon capture and sequestration. Under steady-state conditions, the governing equations (which are coupled PDEs [1]) reduce to a non-linear system of algebraic equations that relate pressures at junctions/nodes with mass flows in the pipes and other devices throughout the network. The solution of such a non-linear system is a crucial ingredient in various decision support tools for system design and operation that aim to ensure safe and reliable transport of fluids through the network.
Non-linear algebraic systems may have no solution, a unique solution or even multiple solutions, and before designing algorithms to solve such systems, it is important to answer the question of existence and uniqueness of solutions if possible. The question of existence is not an abstruse mathematical curiosity, but one which has fundamental significance from the viewpoint of modelling, control as well as computation. The equations representing the model of a real system are a mathematical construct that aims to mimic the physical response of real systems. While the occurence of a physical phenomena may be undeniable and obvious, the verification that the equations can simulate this phenomena with fidelity is tantamount to proving existence of a solution under certain hypotheses or assumptions. Thus, investigation of existence of a solution is linked to the validity of the model itself. In the control of linear systems, the existence of a solution is rarely spoken about since most linear ODE systems under consideration are almost always regular and guarantee existence and uniqueness. However, in case of nonlinear systems, notions such as controllability cannot be well-defined unless it is known that a solution exists [2]. Finally, from the perspective of computation, existence of a solution is an assurance that if successively finer numerical approximations do not converge, it is the algorithm that needs modification and not the problem.
Notwithstanding the connection to these fundamental issues, rarely do systems arising in practice afford us either the luxury of first proving/determining existence of a solution before turning to algorithms for approximating it, or the courtesy of readily satisfying the hypotheses required for the invocation of some mathematical theorem that can assert the existence of a solution. Thus, research efforts march onwards to develop computational algorithms assuming that a solution exists, with uniqueness of solution lending support when applicable. Such is the case even with a well-established, justly celebrated model like the Navier-Stokes equation for incompressible fluids. The subject of this article is the steady-state equations governing adiabatic, unidirectional fluid flow in networks. The canonical form of these equations corresponds to different contexts, such as natural gas delivery, water distribution, or transport of carbon dioxide in carbon capture and sequestration. However, the case of gas flow is considered here since it is the most complicated of the three due to the compressible nature and equation of state that governs the density.
For gas flow, the non-linear operator has been shown to have a unique solution, if one exists [1, 3, 4, 5], and several attempts have been made to develop efficient computational algorithms [6, 7, 4, 1] agnostic to the grid topology and size that can determine a solution. However, none of these algorithms succeed with large grid sizes, and the absence of a definite result on existence of solutions hobbles further development; for it is not possible to distinguish between algorithm failure and non-existence of a solution.
In this letter, we show that a unique solution exists for such non-linear gas flow systems if the term solution is interpreted in terms of potentials and flows rather than pressures and flows. This result applies to other fluid flow networks as well and is a significant improvement from [1], where the authors conclude that even for the simple case of flow in a single pipe, a solution may not exist. A careful examination reveals that their definition of a solution is in terms of pressures and flows and hence does not contradict the result presented here.
We use topological degree theory and homotopy invariance [8, 9] to assert existence of a solution by starting with a linear system (where the degree computation is trivial) that can be continuously transformed to obtain the non-linear system of interest. There are subtle but crucial differences from homotopic continuation methods [10, 11, 12] which are constructive and postulate a trajectory connecting the solution of a known problem to the solution of the non-linear system of interest. Homotopic continuation is mainly used as a numerical algorithm [13, 14], but theoretical assertions require that the system satisfy very stringent hypotheses that are verifiable only in rare cases [15] where the system has some special structure. In contrast, assertions of existence following degree theory and homotopy invariance as used here, require weaker assumptions that even allow the Jacobian to be singular at a solution.
The article is structured as follows: After setting up the notation and recapitulating the governing equations in Section 2, we state and summarize the propositions and assertions required in Section 3 before using those to prove the main result in Section 4 and concluding with Section 5.
2 Mathematical Preliminaries
Let $G(V,E)$ denote the graph of the pipeline network, where $V$, $E$ are the set of junctions and edges of the network respectively. The edge set is given by $E=P\bigcup C$ where $P$ and $C$ are the set of pipes and compressors respectively. A pipe $(i,j)$ in $P$ connects the junctions $i$ and $j$ with a resistance coefficient $\beta_{ij}>0$. Also, a compressor $(i,j)$ in $C$ connects junctions $i$ and $j$, which we assume are geographically co-located. For each junction $i$, we let $p_{i},q_{i}$ denote the nodal value of the pressure and the injection at $i$ respectively. For any edge $(i,j)$ (pipe or compressor) that connects junctions $i$ and $j$, we let $\phi_{ij}$ denote the steady flow through the edge. Finally, for each compressor $(i,j)$, we use $\alpha_{ij}\geqslant 1$ to denote the specified pressure boost ratio or compressor ratio. It is assumed that the direction of flow is in the direction of pressure boost, i.e., $i\to j$, and hence $\phi_{ij}>0$ for a physically reasonable compressor $(i,j)$. Each junction $i$ is assigned to one of two mutually disjoint sets consisting of the slack junctions $N_{s}$ and non-slack junctions $N_{ns}$ respectively, depending on whether pressure $p_{i}$ or gas injection $q_{i}$ at the node is specified. Thus $V=N_{s}\cup N_{ns}$.
There are certain assumptions we must make about the network in order to ensure the validity of the conclusions and assertions that follow. These assumptions are not restrictive, serving instead to eliminate certain impractical but pathological cases, and are stated as follows [1]:
  1. (A1)
    There is at least one slack junction, i.e., $|N_{s}|\geqslant 1$.
  2. (A2)
    When $|N_{s}|\geqslant 2$, a path connecting two slack junctions must consist of at least one pipe.
  3. (A3)
    Any cycle must consist of at least one pipe.
The nature of the gas is prescribed by the equation of state chosen to model the pressure-density relationship of the gas and it determines a potential function $\pi:\mathbb{R}\to\mathbb{R}$ through
[…]
Two important cases of interest are that of the ideal gas and a non-ideal gas that follows the CNGA equation of state [16] for which we have:
$\pi^{\mathrm{ideal}}(p)=p^{2}/2\;,$ (2a)
$\pi^{\mathrm{cnga}}(p)=b_{1}p^{2}/2+b_{2}p^{3}/3\;,$ (2b)
with
[…]
Anticipating subsequent development, we note the following
Lemma 1.
If for all $x,\alpha\in\mathbb{R}$, $\pi(\alpha x)=g(\alpha)\pi(x)$ for a fixed $g:\mathbb{R}\to\mathbb{R}$, then it is necessary and sufficient that $\pi,g$ have the form $\pi(x)=kx^{r},g(\alpha)=\alpha^{r}$ for some $k,r\in\mathbb{R}$. It is assumed that $\pi,g$ are differentiable and non-constant.
Proof.
The sufficiency of the stated representation is obvious. For necessity, observe that if $x=0$, $\pi(0)=g(\alpha)\pi(0)$. If $\alpha=0$, $\pi(0)=g(0)\pi(x)$. These two conditions imply that $\pi(0)=g(0)=0$ in order to rule out the trivial case of constant functions. Now consider $\alpha x\neq 0$. Differentiating with respect to $x$ and $\alpha$, one has $\pi^{\prime}(\alpha x)\alpha=g(\alpha)\pi^{\prime}(x)$ and $\pi^{\prime}(\alpha x)x=g^{\prime}(\alpha)\pi(x)$ respectively. Together, they imply $\alpha g^{\prime}(\alpha)/g(\alpha)=x\pi^{\prime}(x)/\pi(x)=r$ (for some $r\in\mathbb{R}$) from which we may conclude that $\pi(x)=kx^{r},g(\alpha)=c\alpha^{r}$ for some $k,c\in\mathbb{R}$. However, this immediately yields $c=1$. ∎
The potential function $\pi=\pi^{\mathrm{ideal}}$ satisfies Lemma 1 (with $k=1/2,r=2$) while $\pi=\pi^{\mathrm{cnga}}$ does not.
The system of equations for the network may now be summarized as follows [1], with the understanding that quantities with a *** superscript constitute given data:
$\mathcal{G}\mathcal{F}_{p}:$ $\pi(p_{i})-\pi(p_{j})=\beta^{*}_{ij}\phi_{ij}|\phi_{ij}|\quad\forall(i,j)\in P,$ (4a)
$\mathcal{G}\mathcal{F}_{p}:$ $\alpha^{*}_{ij}p_{i}-p_{j}=0\quad\forall(i,j)\in C,$ (4b)
$\mathcal{G}\mathcal{F}_{p}:$ $\underset{(j,i)\in E}{\sum}\phi_{ji}-\underset{(i,j)\in E}{\sum}\phi_{ij}=q^{*}_{i}\;\forall i\in N_{ns},$ (4c)
$\mathcal{G}\mathcal{F}_{p}:$ $p_{i}=p^{*}_{i}\;\forall i\in N_{s}.$ (4d)
Definition 1.
If there exists $\bm{p}\in\mathbb{R}^{|V|},\bm{\phi}\in\mathbb{R}^{|E|}$ that satisfies the $\mathcal{G}\mathcal{F}_{p}$ system (4d), the system is solvable, or alternatively, a generalized pressure-solution to system (4d) exists.
Proposition 1.
Given assumptions (A1)–(A3), and a monotonic potential function $\pi$, solutions to the $\mathcal{G}\mathcal{F}_{p}$ system (4d), if they exist, are unique.
Proof.
See [3, 1]. ∎
If we denote a nodal potential $\pi_{i}=\pi(p_{i})\;\forall i\in V$, we can eliminate the pressures in favour of the potentials in (4a) and (4d). This leaves (4b) as the only remaining equation where variables corresponding to nodal pressures are present.
The equation (4b) relating pressures $p_{i},p_{j}$ implies that $\pi(\alpha^{*}_{ij}p_{i})=\pi_{j}$. When the potential function $\pi$ is as in Lemma 1, we obtain an equation relating potentials $\pi_{i},\pi_{j}$ for some suitable $\gamma^{*}_{ij}=g(\alpha^{*}_{ij})$. Thus the system (4d) is transformed into a new system:
$\mathcal{G}\mathcal{F}_{\pi}:$ $\pi_{i}-\pi_{j}=\beta^{*}_{ij}\phi_{ij}|\phi_{ij}|\quad\forall(i,j)\in P,$ (5a)
$\mathcal{G}\mathcal{F}_{\pi}:$ $\gamma^{*}_{ij}\pi_{i}-\pi_{j}=0\quad\forall(i,j)\in C,$ (5b)
$\mathcal{G}\mathcal{F}_{\pi}:$ $\underset{(j,i)\in E}{\sum}\phi_{ji}-\underset{(i,j)\in E}{\sum}\phi_{ij}=q^{*}_{i}\;\forall i\in N_{ns},$ (5c)
$\mathcal{G}\mathcal{F}_{\pi}:$ $\pi_{i}=\pi^{*}_{i}\;\forall i\in N_{s}.$ (5d)
Remark 1.
Clearly, transforming $\mathcal{G}\mathcal{F}_{p}$ system (4d) to $\mathcal{G}\mathcal{F}_{\pi}$ system (5d) is possible when the ideal gas equation of state is assumed but the CNGA equation of state forbids such a transformation in (4b). Further analysis of the system corresponding to the CNGA equation of state is postponed to Sec. 4.
Definition 2.
If there exists $\bm{\pi}\in\mathbb{R}^{|V|},\bm{\phi}\in\mathbb{R}^{|E|}$ that satisfies the $\mathcal{G}\mathcal{F}_{\pi}$ system (5d), the system is solvable, or alternatively, a generalized potential-solution to system (5d) exists.
Lemma 2.
If the $\mathcal{G}\mathcal{F}_{\pi}$ system (5d) is solvable, then the solution is unique.
Proof.
One can interpret the $\mathcal{G}\mathcal{F}_{\pi}$ system (5d) as a special case of $\mathcal{G}\mathcal{F}_{p}$ system (4d) with the equation of state $\rho(p)=1$, and $\pi(p)=p$. By Proposition 1, solutions to the $\mathcal{G}\mathcal{F}_{p}$ system (4d), if they exist, are unique. Then it is clear that if the $\mathcal{G}\mathcal{F}_{\pi}$ system (5d) is solvable, then the solution is unique. ∎
For given Definitions 1 and 2 assuming Lemma 1, note that solvability of $\mathcal{G}\mathcal{F}_{p}$ system (4d) implies solvability of $\mathcal{G}\mathcal{F}_{\pi}$ system (5d). However, the converse is false, for even if system (5d) is solvable, for some $\pi_{i}\in\mathbb{R}$, it is possible that there is no $p_{i}\in\mathbb{R}$ such that $\pi(p_{i})=\pi_{i}$. In any case, the equation (5b) will pose no problems since the form of $\pi$ is dictated by Lemma 1.
Remark 2.
We do not impose any restrictions (such as sign or domain) motivated by physical considerations on $\bm{\pi}$, $\bm{p}$ or $\bm{\phi}$ in Definitions 1 and 2, hence we use the term generalized solution.
The questions we would like to answer are the following:
  1. (Q1)
    Do the systems (4d) or (5d) for an ideal gas have a solution?
  2. (Q2)
    For a non-ideal gas (with the CNGA equation of state), does the system (4d) have a solution?
  3. (Q3)
    Can we prescribe a domain within which solutions (if they exist) must lie?
We can answer (Q3) for the system (5d) by estimating the upper bound using the given data as follows:
Proposition 2.
Given a network $G(V,E)$, with sets $P,C$, $N_{ns},N_{s}$ and data as in system (5d), we define the following quantities:
$\beta_{m}\triangleq\max_{(i,j)\in P}\beta^{*}_{ij},\quad\alpha_{m}\triangleq\max_{(i,j)\in C}\alpha^{*}_{ij},$
$\phi_{m}\triangleq\left(\max_{i\in N_{ns}}|q^{*}_{i}|\right)\cdot|N_{ns}|,$
$\pi_{m}\triangleq\left(\max_{i\in N_{s}}|\pi^{*}_{i}|+|P|\beta_{m}\phi_{m}^{2}\right)\cdot(\alpha_{m})^{|C|},$
$L\triangleq\max(\phi_{m},\pi_{m}).$
Then the hypercube
$\mathcal{D}_{\mathcal{H}}=\{\bm{x}\in\mathbb{R}^{|V|+|E|}:|x_{i}|<L\}$ (6)
is an open, bounded set that contains all solutions (if they exist) of system (5d).
The upper bounds above have been derived by considering the maximum possible flow into any single pipe, then calculating the maximum possible pressure by considering a node connected to the slack node with maximum pressure, and furthermore, taking the maximal possible compression achievable in the full network.
In Section 4, we shall find that the answer to (Q1) is that for an ideal gas, $\mathcal{G}\mathcal{F}_{\pi}$ system (5d) has a unique generalized potential solution. If the solution admits non-negative potentials, then the $\mathcal{G}\mathcal{F}_{p}$ system (4d) is also solvable, having at least one generalized pressure solution.
3 Background
Before proving the existence results outlined in the previous section, we state (without proof) some mathematical results that we shall soon invoke.
Definition 3.
A point $\bm{y}$ is a regular value of a map $\bm{f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ if $\det(D\bm{f}(\bm{x}))\neq 0\;\forall\;\bm{x}\in\bm{f}^{-1}\{\bm{y}\}$.
Definition 4.
Given a map $\bm{f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and an open bounded domain $\mathcal{D}\subset\mathbb{R}^{n}$ such that
  1. 1.
    $\bm{f}$ is continuously differentiable on $\mathcal{D}$,
  2. 2.
    $\bm{y}$ is a regular value of $\bm{f}$,
  3. 3.
    $\bm{y}\notin\bm{f}(\partial\mathcal{D})$,
the degree [8, Chapter 6] of $\bm{f}$ at $\bm{y}$ given $\mathcal{D}$ is defined as
$\mathrm{deg}(\bm{f},\mathcal{D},\bm{y})=\sum_{i}\mathrm{sgn}(\det(D\bm{f}(\bm{x}_{i})))\;\forall\;\bm{x}_{i}\in\bm{f}^{-1}\{\bm{y}\}.$
If $\bm{f}^{-1}\{\bm{y}\}$ is empty we define $\mathrm{deg}(\bm{f},\mathcal{D},\bm{y})=0$.
Remark 3.
Given a domain $\mathcal{D}\subset\mathbb{R}^{n}$ and an invertible linear operator $\bm{A}$, $\mathrm{deg}(\bm{A},\mathcal{D},\bm{y})\neq 0$ for any $\bm{y}\in\mathcal{D}$.
The definition of degree can be extended to allow non-regular values without affecting subsequent results, namely the following two propositions. The proofs and other details of degree theory may be found in [8, Chapter 6] or [9].
Proposition 3.
If $\mathrm{deg}(\bm{f},\mathcal{D},\bm{0})\neq 0$, then $\bm{f}$ must have a zero/root in $\mathcal{D}$.
Proposition 4 (Homotopy invariance).
Let $\mathcal{D}\subset\mathbb{R}^{n}$ be open and bounded and let $H:\mathrm{cl}({\mathcal{D}})\times[0,1]\to\mathbb{R}^{n}$ be a continuous map on $\mathrm{cl}({\mathcal{D}})\times[0,1]$ where $\mathrm{cl}({\mathcal{D}})$ is the closure of the set $\mathcal{D}$. Suppose further that $H(\bm{x},s)\neq\bm{0}$ in $\partial\mathcal{D}\times[0,1]$. Then $\mathrm{deg}(H(\bm{x},s),\mathcal{D},\bm{0})$ is constant for all $s\in[0,1]$. Thus if $\mathrm{deg}(H(\bm{x},0),\mathcal{D},\bm{0})\neq 0$, then $H(\bm{x},s)=\bm{0}$ has a solution in $\mathcal{D}$ for every $s\in[0,1]$.
In order to use Proposition 4, first we recognize that the solution of system  (4d) and (5d) may be thought of as a zero or root of a non-linear operator on $\mathbb{R}^{|E|+|V|}$. In order to conform to Proposition 4, we need to modify system (4d) or (5d) to introduce a scalar parameter $s\in[0,1]$ such that the operator has the form $F:\mathbb{R}^{|E|+|V|}\times\mathbb{R}\longrightarrow\mathbb{R}^{|E|+|V|}$. This is easily accomplished in system (7d) so that setting $s=1$ recovers system (5d) as follows:
$F((\bm{\phi},\bm{\pi}),s):$ $\pi_{i}-\pi_{j}=\beta^{*}_{ij}\phi_{ij}|\phi_{ij}|^{s}\quad\forall(i,j)\in P,$ (7a)
$F((\bm{\phi},\bm{\pi}),s):$ $\gamma^{*}_{ij}\pi_{i}-\pi_{j}=0\quad\forall(i,j)\in C,$ (7b)
$F((\bm{\phi},\bm{\pi}),s):$ $\underset{(j,i)\in E}{\sum}\phi_{ji}-\underset{(i,j)\in E}{\sum}\phi_{ij}=q^{*}_{i}\;\forall i\in N_{ns},$ (7c)
$F((\bm{\phi},\bm{\pi}),s):$ $\pi_{i}=\pi^{*}_{i}\;\forall i\in N_{s}.$ (7d)
Remark 4.
$F((\bm{\phi},\bm{\pi}),s)$ in (7d) is once-differentiable with a continuous derivative.
The following result (restated with slight modification from [1]) that guarantees the invertibility of a linear system with particular structure is essential for the subsequent analysis.
Proposition 5.
The only solution to the following homogeneous linear system in ($(\{x_{k}\},\{y_{ij}\})$, where $k\in V$, $(i,j)\in E$ for an admissible gas network is $x_{k}=y_{ij}=0$, where the real numbers $a^{*}_{i}\neq 0$, all $a^{*}_{i}$ have the same sign, and for every cycle $\mathfrak{C}\in G(V,E)$, there exists at least one $(i,j)\in\mathfrak{C}$ such that $b^{*}_{ij}\neq 0$:
$a^{*}_{i}x_{i}-a^{*}_{j}x_{j}-b^{*}_{ij}y_{ij}=0\quad\forall(i,j)\in E,$ (8a)
$\underset{(j,i)\in E}{\sum}y_{ji}-\underset{(i,j)\in E}{\sum}y_{ij}=0\;\forall i\in N_{ns},$ (8b)
$x_{i}=0\quad\forall i\;\in N_{s}.$ (8c)
4 Results and Implications
Theorem 1.
A unique solution to $\mathcal{G}\mathcal{F}_{\pi}$ system (5d) exists.
Proof.
We shall appeal to Proposition 4 to assert the existence of a solution. Accordingly, we first observe that $F((\bm{\phi},\bm{\pi}),0)$ is a linear system of the same form as Proposition 5 and hence has a unique solution in the open hypercube $\mathcal{D}_{\mathcal{H}}$ in (6) constructed in Proposition 2. Moreover, $F((\bm{\phi},\bm{\pi}),s)$ is continuous on $\mathrm{cl}(\mathcal{D}_{\mathcal{H}})$ and by construction no solution, if it exists, can lie on the boundary of the hypercube. Hence by Proposition 4, $F((\bm{\phi},\bm{\pi}),s)$ has a solution in $\mathcal{D}_{\mathcal{H}}$ for every $s\in[0,1]$. Thus, for $s=1$, we conclude that the system (5d) has a generalized-potential solution. Once existence is assured, uniqueness follows from Lemma 2. ∎
The following corollary follows easily from Theorem 1:
Corollary 1.
If the (unique) solution to (5d) has values of the potential $\pi_{i}\in\mathbb{R}\;\forall i\in V$ such that they can be inverted to find pressures $p_{i}\in\mathbb{R}$, such that $\pi(p_{i})=\pi_{i}$, then a generalized pressure solution exists for system (4d).
We note in passing that for $\pi^{\mathrm{ideal}}$ (2a), inversion is possible only for non-negative values but since the expression for $\pi^{\mathrm{cnga}}$ (2b) is a cubic polynomial, a preimage can be found corresponding to any real value of the potential.
4.1 Ideal Gas
Theorem 1 and Corollary 1 have far reaching implications, laying to rest all ambiguity about the solvability of the steady-state gas flow equations on networks for the case of the ideal gas. They assure us that steady-state network flow for an ideal gas (corresponding to system (5d)) always has a unique, generalized potential solution. Furthermore, the expression for $\pi$ in (2a) makes it clear that a generalized-pressure solution can be derived contingent on non-negative potentials.
It was pointed out in Remark 2 that our definition of generalized solution allowed the potentials, pressures and flows to take any real values. However, at this juncture, we can allow physical considerations to make a distinction between the equations being solvable and feasible.
As outlined in [1], for both the ideal and the non-ideal gas (with CNGA equation of state), a feasible solution will have non-negative pressures and non-negative flow in compressors. Otherwise, the solution/system is infeasible.
In light of this statement, one may qualify the questions (Q1), (Q2) asked earlier to now read:
  1. (Q4)
    Do the systems (4d) or (5d) for an ideal gas have a feasible solution?
  2. (Q5)
    For a non-ideal gas (with the CNGA equation of state), does the system (4d) have a feasible solution?
From the theory developed thus far, one may definitively conclude apropos of (Q4) that the system (5d) for the ideal gas is infeasible if any component of the generalized potential solution or compressor flow is negative, but feasible otherwise.
4.2 Non-ideal Gas
(a) $g(\alpha)=\gamma_{\mathrm{opt}}$
(b) $g(\alpha)=\alpha^{2}$
Figure 1: The plots above show the magnitude and distribution of the relative error $\mathrm{e}_{\mathrm{rel}}(\alpha,p)$ in (10) for two different choices of $g(\alpha)$. The colour bars demonstrate how the error is reduced drastically by choosing the optimal $g(\alpha)$ in (9). With this choice, we may write $\pi(\alpha p)\approx\gamma_{\mathrm{opt}}\pi(p)$ so that the system (4d) for a non-ideal gas is approximated by (5d).
We have thus far made no assertions about the existence of steady-state solutions when the gas obeys the CNGA equation of state and thus no attempt to answer (Q2) and (Q5) articulated earlier. For the existence of solutions to system (4d) for a non-ideal gas, it is tempting to construct a continuous homotopy function $\chi(p,s)$ that satisfies $\chi(p,0)=\pi^{\mathrm{ideal}}(p)$ and $\chi(p,1)=\pi^{\mathrm{cnga}}(p)$ to replace $\pi(p)$ in system (4d) and reuse the artifice of Proposition 4 to analyse the system (4d). However, it is not clear how to estimate an upper bound for the solution due to the term $\chi(p,s)$ analogous to Proposition 2; indeed we cannot be sure that no solution will lie on the prescribed boundary for any $s\in[0,1]$.
In abandoning this approach and conceding the futility of the task associated with (Q2) and (Q5), it is natural to consider if the theory developed for the ideal gas could prove useful in some other way for the non-ideal gas. Towards this end, we look to examine the possibility of approximating system (4d) by (5d) for the CNGA equation of state. We already know that because the potential function for the CNGA equation of state (2b) violates Lemma 1, (5b) does not follow from (4b).
However, if we could write $\pi(\alpha p)\approx g(\alpha)\pi(p)$ for some function $g$, then (5d) could be a system that approximates (4d) for the CNGA equation of state in the sense that the solution to (5d) is an acceptable approximant to the solution of (4d) when the latter exists. To this end, we could look to determine
$g(\alpha)\in\mathrm{span}\{1,\alpha,\alpha^{2},\alpha^{3}\}$
such that the residual $\pi(\alpha p)-g(\alpha)\pi(p)$ is minimized in a least-squares sense implied by the quantity
$\int_{\alpha_{\mathrm{min}}}^{\alpha_{\mathrm{max}}}\int_{p_{\mathrm{min}}}^{p_{\mathrm{max}}}\left(\pi(\alpha p)-g(\alpha)\pi(p)\right)^{2}\mathrm{d}p\;\mathrm{d}\alpha.$
[…]
$\gamma_{\mathrm{opt}}=g(\alpha)\approx 0.9\alpha^{2}+0.1\alpha^{3}.$ (9)
The significance of this choice can be evaluated by comparing the magnitude and the distribution of the relative error
$\mathrm{e}_{\mathrm{rel}}(\alpha,p)=\left|\dfrac{\pi(\alpha p)-g(\alpha)\pi(p)}{\pi(\alpha p)}\right|$ (10)
for $g(\alpha)=\gamma_{\mathrm{opt}}$ with that of $g(\alpha)=\alpha^{2}$ (which would be applicable for the ideal gas (2a)) in Figure 1. It can be seen that the error is reduced drastically by choosing the optimal $g(\alpha)$ in (9). With this choice, we may write $\pi(\alpha p)\approx\gamma_{\mathrm{opt}}\pi(p)$ so that the system (4d) for a non-ideal gas is approximated by (5d).
Let $\bm{x}_{a}$ be the pressure-flux pair obtained from inversion of the potentials in the solution to system (5d) with the choice $\gamma_{\mathrm{opt}}$ in (5b). Note that $\bm{x}_{a}$ can always be found since a unique generalized potential solution to system (5d) always exists, and a pressure corresponding to a potential can always be found for the cubic polynomial (2b). Suppose that $\bm{x}$ is the solution to the original system (4d). If we denote by $\bm{R}(\bm{y})$ the vector of residuals of the system (4d) evaluated at $\bm{y}$, then $||\bm{R}(\bm{x})||_{\infty}\approx 0$.
Motivated by standard ideas in inversion of linear systems [17], we want an approximate solution to have a small residual, and also be close to the true solution. Thus, the two quantities that measure the acceptability of $\bm{x}_{a}$ as an approximation of $\bm{x}$ are
$\dfrac{||\bm{x}_{a}-\bm{x}||_{\infty}}{||\bm{x}||_{\infty}}\text{ and }||\bm{R}(\bm{x}_{a})||_{\infty}.$
Unfortunately, we are unable to estimate these quantities theoretically. The only qualitative assertion we can make is that the solution $\bm{x}_{a}$ ought to satisfy (4a), (4d) and that the quantity $||\bm{R}(\bm{x}_{a})||_{\infty}$ will be determined by (4b) alone. We also expect that a good approximate solution will serve as a good initial guess in a Newton-Raphson algorithm and lead to convergence within a handful of iterations.
Figure 2: Schematic of the GasLib-40 network. The junctions that connect compressors are shown in red and the slack junction is shown in blue.
[…]
While no claim can be made regarding the existence of solutions to system (4d) in general, these numerical experiments support our assertion that the solution $\bm{x}_{a}$ of system (5d) is a good approximation to the computed solution $\bm{x}$ of system (4d). Moreover, starting with the initial guess $\bm{x}_{a}$, the convergence to $\bm{x}$ was attained in exactly 2 Newton iterations in each instance.
The comparison between the solutions for an ideal gas and a non-ideal gas has already been analyzed in [1] and is hence not considered here.
Figure 3: Cumulative distribution function for GasLib-40 showing the residual $O(10^{-2})$ as well as the relative error $O(10^{-3})$ of the approximate solution for the non-ideal gas.
4.3 Other fluids
The results obtained so far can be generalized to other fluid flow networks once the equation of state that describes the pressure-density relationship is specified. For instance, in water distribution networks or in carbon capture and sequestration, the incompressibility of water and carbon-dioxide implies that the equation of state is $\rho(p)=\rho^{*}$ (a constant), which makes the potential function linear (and simpler) [19] in (1). The other difference is the use of pumping devices which provide a given additive increase in potential or pressure and simplify the analysis [20] in comparison to a given multiplicative increase as in (4b) or (5b) considered for compressors.
5 Conclusion
The arguments and proofs presented in this letter lay to rest the question of existence of solutions to potential-driven, steady-state fluid flow in a network. As the question of existence of solutions had in it the embedded complexity of the network topology and the equation of state of the gas, the results are also applicable to other networks carrying fluids such as water [19] or carbon-dioxide. Importantly, it gives support to computational work by enabling correct diagnosis of algorithmic failure, problem stiffness and non-convergence by answering the question of existence of solutions. Future work could look towards determining solutions by homotopic continuation methods [12] starting with a less complicated system with a guaranteed solution as exemplified by [15].